OpenStudy (nanopoison13):

If object B is 6 times smaller than object A in all dimensions, then how many times is object A larger than object B?

OpenStudy (kittiwitti1):

\[\beta\text{ is }x \text{ times smaller than }\alpha\]\[\beta=\frac{\alpha}{x}\]In this problem it tells you \(x=6\), so input values as specified.

OpenStudy (wolf1728):

However, imagine that object B is cube-shaped that is 6 inches by 6 inches by 6 inches. If object A is 6 times smaller than object B in EVERY direction, then object A should measure 1 inch by 1 inch by 1 inch. Object B's volume would be 6*6*6 = 216 cubic inches Object A's volume would be 1*1*1 = 1 cubic inch So object A's volume would be 216 times smaller than object B.

OpenStudy (snowflakelove15):

It would be 6 times because if object b is 6 times smaller and there asking you how much bigger object a is then you would just change it from smaller to larger.